# From LLMs to LLM-based Agents for Software Engineering: A Survey of Current, Challenges and Future

### v1

@@ I Introduction
Software engineering (SE) has seen its booming research and development with the aid of artificial intelligence techniques. Traditional approaches leveraging neural networks and machine learning have facilitated various SE topics such as bug detection, code synthesis, and requirements analysis [1][2]. However, they often present limitations, including the need for exclusive feature engineering, scalability issues, and the adaptability across diverse codebases. The rise of Large Language Models (LLMs) has embarked on new solutions and findings in this landscape. LLMs, such as GPT [3] and Codex [4], have demonstrated remarkable capabilities in handling downstream tasks in SE, including code generation, debugging, and documentation. These models leverage vast amounts of training data to generate human-like text, offering unprecedented levels of fluency and coherence. Studies have shown that LLMs can enhance productivity in software projects by providing intelligent code suggestions, automating repetitive tasks, even generating entire code snippets from natural language descriptions [5].
@@
To address these challenges, LLM-based agents have emerged [9][10], combining the strengths of LLMs with external tools and resources to enable more dynamic and autonomous operations. These agents leverage recent advancements in AI, such as Retrieval-Augmented Generation (RAG) and tool utilization, to perform more complex and contextually aware tasks [11]. For instance, OpenAI’s Codex has been integrated into GitHub Copilot [12], enabling real-time code suggestions and completion within development environments. Unlike static LLMs, LLM-based agents can perform a wide range of tasks, such as autonomously debugging code by identifying and fixing errors, proactively refactoring code to enhance efficiency or readability, and generating adaptive test cases that evolve alongside the codebase. These features make LLM-based agents a powerful tool for SE, capable of handling more complex and dynamic workflows than traditional LLMs.
Historically, AI agents focused on autonomous actions based on predefined rules or learning from interactions  [13][14]. The integration of LLMs has presented new opportunities in this area, providing the language understanding and generative capabilities needed for more sophisticated agent behaviors. [10] shows that LLM-based agents are capable of autonomous reasoning and decision-making, achieving the third and fourth levels of WS (World Scope) [15], which outlines the progression from natural language processing (NLP) to general AI. In software engineering, LLM-based agents show promise in areas such as autonomous debugging, code refactoring, and adaptive test generation, demonstrating capabilities that approach artificial general intelligence(AGI).
Figure 1: Paper Number for LLMs and LLM-based Agent between 2020-2024
@@
  • RQ1: What are the state-of-the-art techniques and practices in LLMs and LLM-based agents for SE? (Section. IVIX)
  • RQ1: What are the key differences in task performance between LLMs and LLM-based agents in SE applications? (Section. IVIX)
  • RQ2: Which benchmark datasets and evaluation metrics are most commonly used for assessing the performance of LLMs and LLM-based agents in SE tasks? (Section. IVIX and Section. X)
  • RQ3: What are the predominant experimental models and methodologies employed when utilizing LLMs in SE? (Section. X)
II EXISTING WORKS AND THE SURVEY STRUCTURE
@@ II-A Existing works
In recent years, large language models have been primarily applied to help programmers generate code and fix bugs. These models understand and complete code or text based on the user’s input, leveraging their training data and reasoning capabilities. In previous survey papers, such as Angela Fan’s research [8], there has not been much elaboration on requirement engineering. As mentioned in the paper, software engineers are generally reluctant to rely on LLMs for higher-level design goals. However, with LLMs achieving remarkable improvements in contextual analysis and reasoning abilities through various methods like prompt engineering and Chain-of-Thought (COT) [16], their applications in requirement engineering are gradually increasing. Table I summarizes and categorizes the tasks in requirement engineering. Many studies utilize models for requirement classification and generation. Since the collection primarily focuses on the latter half of 2023 and before April 2024, and some papers address multiple tasks, the table does not reflect the exact number of papers we have collected.
@@
39
Figure 2: Paper Distribution
While other works have surveyed LLMs applications in some SE tasks [17] [8] [18], they lack a wider coverage of the general SE area to incorporate recent research developments. More importantly, a focus of LLMs is the main contributions of these works, but there is no distinguish the capabilities between LLMs and LLM-based agents. We summarize the difference between our work and others in Table II, this survey addresses these limitations by distinctly analyzing LLMs and LLM-based agents applications across six SE domains, providing a thorough and up-to-date review. From previous research, it is evident that the performance of LLMs in various applications and tasks heavily depends on the model’s inherent capabilities  [10]. More importantly, earlier surveys often present findings from papers spanning in a wide range of publication dates, leading to significant content disparities for LLMs in different SE tasks. For instance, research in requirement engineering was relatively nascent, resulting in sparse content in this area in previous surveys. The recent rise of LLM-based agents, with their enhanced capabilities and autonomy, fills these gaps. By focusing on the latest researches and clearly differentiating between LLMs and LLM-based agents, our survey provides a thorough and in-depth overview of how these technologies are applied and new opportunities they bring to SE.
In summary, we have collected a total of 117 papers directly relevant to this topic, covering the six SE domains mentioned earlier as shown in Figure.1. Our analysis distinguishes between LLM and LLM-based agent contributions, offering a comparative overview and addressing the limitations of previous surveys. Considering the novelty nature of the LLM-based agents field and the lack of standardized benchmarks, this work seeks to offer a detailed review that can guide future research and provide a clearer view of the potential of these technologies in SE.
TABLE II: Comparison between Our Work and the Existing Work for LLM in SE
@@ II-B Methodology
The paper collection process primarily involved searching DBLP and arXiv databases, focusing on recent studies from the latter half of 2023 to May 2024. This approach ensured the inclusion of the latest research. We filtered out non-LLM-related papers and those with fewer than seven pages. To further refine our selection, we used keywords in Table III to search SE-related works. We then manually screened the remaining papers to remove any with formatting errors or student projects. Additionally, we employed a snowballing search technique to capture significant works that might have been missed initially. Overall, we identified 117 relevant papers. Figure. 2 presents the distribution of these papers across the six SE domains and the proportion of LLM-based agents studies. However, some papers can be counted as multi-class fields so the literature review in the figure is more than 117 in total.
@@ II-C Overall Structure of the Work
The remainder of this paper is organized as follows Section 2 Introduces the architectures and background of LLMs and LLM-based agents, including an overview of RAG, tool utilization, and their implications for SE. Section 3-8 is the comparative analysis which Summarizes and compares the datasets, tasks, benchmarks, and metrics used in LLM and LLM-based agents studies across the six SE domains. Section 9 is the general discussion, section 10 is the final conclusion.
TABLE III: Keywords for Software Engineering Topics
@@ III-A Large Language Model
There is an inherent connection between large language models and natural language processing (NLP), with the historical development of natural language technologies tracing back to the 1950s. The earliest attempts to generate language dialogues through machines using specific rules can be traced to the period between 1950 and 1970. The advent of machine learning technologies in the 1980s and the groundbreaking introduction of neural networks in the 1990s indicated a new era for NLP  [23]. These advancements facilitated significant progress in the NLP field, especially in the development of technologies for text translation and generation. The development of Long Short-Term Memory (LSTM) and Recurrent Neural Networks (RNN) during this period enabled more effective handling of the sequential nature of language data [24] [25]. These models addressed challenges associated with the lack of dependency in context, thereby enhancing the application of NLP in various domains.
@@ III-C Large Language Model Based Agent
Figure 3: Illustration of Common Data Augmentation Methods
Nowadays, various agent systems are emerging and they relies on large language models to make judgments, combined with techniques such as few-shot learning and multi-turn dialogue for model fine-tuning. However, due to the lack of datasets and the novelty of LLM-based agents, many researchers employ different methods for data augmentation. Common approaches include synonym replacement, where words in the text are replaced with synonyms from the same domain to increase textual diversity; back-translation, where the text is translated into another language and then back to the original language to generate new texts with slightly different grammatical structures and word choices; Paraphrasing refers to the new dialogue which similar in context but slightly different in expression created through manual or automated means; Synthetic Data Generation refers to use pretrained model to make a synthetic data generation as shown in Figure.3. In 2023, Chenxi Whitehouse explored using LLMs for data augmentation to enhance the performance of multilingual commonsense reasoning datasets, especially under conditions of extremely limited training data [39]. The study employed various LLMs (such as Dolly-v2, StableVicuna, ChatGPT, and GPT-4) to generate new data. These models were prompted to create new examples similar to the original data, thereby increasing the diversity and quantity of training data. Prompt engineering was mentioned as an essential skill for effectively interacting with LLMs. By applying these prompt patterns, users can efficiently customize their dialogues with LLMs ensuring the generation of high-quality outputs and achieving complex automated tasks. In 2023, Jules White introduced a set of methods and patterns to enhance prompt engineering, optimizing interactions with LLMs such as ChatGPT [40]. This study categorizes prompt engineering into five main areas: Input Semantics, Output Customization, Error Identification, Prompt Improvement, and Interaction, to address a wide range of problems and adapt to different fields.
@@ III-D Single and Multi-Agents
Multi-agent systems involve the collaboration of multiple LLMs or agents to tackle complex tasks effectively. These systems fully utilize the advantages of multiple models, with each model specializing in specific aspects of the task to reduce overhead caused by multi-processes in single agents, the collaboration among agents allows for more sophisticated and robust problem-solving capabilities. Due to their exceptional capabilities, more researchers are beginning to explore the field of Multi-LLM based agents and start applying into software engineering domains. In 2024, a lot of researchers adopt the multi-agent system into the practical experiments [43] [44].
@@ III-E LLM in Software Engineering
Currently, there is no comprehensive and accurate definition of the capabilities an LLM must exhibit to be considered an llm-based agent. Since the application of LLMs in software engineering is relatively broad and some frameworks already behave certain levels of autonomy and intelligent, this study defines the distinction between LLM and LLM-based agents based on mainstream definitions and literature from the first half of 2024. In this survey, an LLM architecture can be called an agent when it satisfy the Table IV.
TABLE IV: Criteria for LLM-based agent
@@ IV-A LLMs Tasks
In the field of requirement engineering, LLMs have demonstrated significant potential in automating and enhancing tasks such as requirement elicitation, classification, generation, specification generation, and quality assessment. Requirement classification and extraction is a crucial task in requirement engineering during the development process. It is common to encounter situations where clients present multiple requirements at once, necessitating manual classification by developers. By categorizing requirements into functional and non-functional requirements, developer can better understand and manage them, thanks to the strong performance of LLMs in classification tasks, many relevant frameworks have been developed. The PRCBERT framework, utilizing the BERT pre-trained language model, transforms classification problems into a series of binary classification tasks through flexible prompt templates, significantly improving classification performance [48]. Studies have shown that the PRCBERT achieved an F1 score of 96.13% on the PROMISE dataset which outperform the previous state-of-arts NoRBERT [49] and BERT-MLM models [31]. Additionally, the application of ChatGPT in requirement information retrieval has shown promising results, by classifying and extracting information from requirement documents, ChatGPT achieved comparable or even better $F\beta$ scores under zero-shot settings, particularly in feature extraction tasks, where its performance surpassed baseline models [50]. As seen in Table I, there is also substantial literature and research on using LLMs to automatically generate requirements and descriptions in requirement engineering.
@@
By automating the generation and description of requirements, the efficiency and accuracy of requirement elicitation can be improved. Research indicates that LLMs hold significant potential in requirements generation task. For example, using ChatGPT to generate and gather user requirements, studies found that participants with professional knowledge could use ChatGPT more effectively, indicating the influence of domain expertise on the effectiveness of LLM-assisted requirement elicitation [51]. The study employed qualitative assessments of the LLMs’ output against predefined criteria for requirements matches, including full matches, partial matches, and the relevancy of the elicited requirements, although their success varied depending on the complexity of the task and the experience of the users, the result showing that LLMs could effectively assist in eliciting requirements, and its particularly useful in identifying, and suggesting requirements based on the large corpus of training data they provided. The SRS (Software Requirement Specification) generation is an important task which the developer normally spent a lot of time to refine and verified. In [52], researchers use both iterative prompting and a single comprehensive prompt to assess the performance of LLMs to generate SRS. The experiment conducted on GPT-4 and CodeLlama-34b one close-source LLM and one open-source LLM for comprehensive evaluation, the generated SRS will compare with human-crafted SRS and finally scored by the likert scale. The result indicate that, the human-generated SRS was overall superior, but CodeLlama often came close, sometimes outperforming in specific categories. The CodeLlama scored higher in completeness and internal consistency than GPT-4 but less concise, so this stuy demonstrated the potential of using fine-tuned LLMs to generate SRS and increase the overall project productivity. Another paper also explores using LLMs for generating specifications. In [53], the authors introduce a framework called SpecGen for generating program specifications. The framework primarily uses GPT-3.5-turbo as the base model and employs prompt engineering combined with multi-turn dialogues to generate the specifications. SpecGen applies four mutation operators to modify these specifications and finally uses a heuristic selection strategy to choose the optimal variant. The results show that SpecGen can generate 70% of the program specifications, outperforming traditional tools like Houdini [54] and Daikon [^1] https://github.com/codespecs/daikon.
@@
Furthermore, designing prompt patterns can significantly enhance LLMs’ capabilities in tasks such as requirement elicitation and system design. The paper provides a catalog of 13 prompt patterns, each aimed at addressing specific challenges in software development [55]. The experiments test the efficacy of these patterns in real world scenarios to validate their usefulness. By applying different prompt patterns, the study found that these patterns could help generate more structured and modular results and reduce common errors. Automated requirement completeness enhancement is another important benefit brought by the LLMs in requirement generation. The study [56] use BERT’s Masked Language Model (MLM) can detect and fill in missing parts in natural language requirements, significantly improving the completeness of requirements. BERT’s MLM achieved a precision of 82%, indicating that 82% of the predicted missing terms were correct.

### v2

@@ I Introduction
Software engineering (SE) has seen its booming research and development with the aid of artificial intelligence techniques. Traditional approaches leveraging neural networks and machine learning have facilitated various SE topics such as bug detection, code synthesis, and requirements analysis [1, 2]. However, they often present limitations, including the need for exclusive feature engineering, scalability issues, and the adaptability across diverse codebases. The rise of Large Language Models (LLMs) has embarked on new solutions and findings in this landscape. LLMs, such as GPT [3] and Codex [4], have demonstrated remarkable capabilities in handling downstream tasks in SE, including code generation, debugging, and documentation. These models leverage vast amounts of training data to generate human-like text, offering unprecedented levels of fluency and coherence. Studies have shown that LLMs can enhance productivity in software projects by providing intelligent code suggestions, automating repetitive tasks, even generating entire code snippets from natural language descriptions [5].
@@
To address these challenges, LLM-based agents have emerged [9, 10], combining the strengths of LLMs with external tools and resources to enable more dynamic and autonomous operations. These agents leverage recent advancements in AI, such as Retrieval-Augmented Generation (RAG) and tool utilization, to perform more complex and contextually aware tasks [11]. For instance, OpenAI’s Codex has been integrated into GitHub Copilot [12], enabling real-time code suggestions and completion within development environments. Unlike static LLMs, LLM-based agents can perform a wide range of tasks, such as autonomously debugging code by identifying and fixing errors, proactively refactoring code to enhance efficiency or readability, and generating adaptive test cases that evolve alongside the codebase. These features make LLM-based agents a powerful tool for SE, capable of handling more complex and dynamic workflows than traditional LLMs.
Historically, AI agents focused on autonomous actions based on predefined rules or learning from interactions  [13, 14]. The integration of LLMs has presented new opportunities in this area, providing the language understanding and generative capabilities needed for more sophisticated agent behaviors. [10] shows that LLM-based agents are capable of autonomous reasoning and decision-making, achieving the third and fourth levels of WS (World Scope) [15], which outlines the progression from natural language processing (NLP) to general AI. In software engineering, LLM-based agents show promise in areas such as autonomous debugging, code refactoring, and adaptive test generation, demonstrating capabilities that approach artificial general intelligence(AGI).
Figure 1: Number of Papers Collected on LLMs and LLM-based Agents from 2020 to 2024.
@@
  • RQ1: What are the state-of-the-art techniques and practices in LLMs and LLM-based agents for SE? (Section IVIX)
  • RQ2: What are the key differences in task performance between LLMs and LLM-based agents in SE applications? (Sections IVIX)
  • RQ3: Which benchmark datasets and evaluation metrics are most commonly used for assessing the performance of LLMs and LLM-based agents in SE tasks? (Section IVIX and Section X)
  • RQ4: What are the predominant experimental models employed when utilizing LLMs in SE? (Section X)
II EXISTING Surveys AND THE SURVEY STRUCTURE
@@ II-A Existing Works
In recent years, large language models have been primarily applied to help programmers generate code and fix bugs. These models understand and complete code or text based on the user’s input, leveraging their training data and reasoning capabilities. In previous survey papers, such as Angela Fan’s research [8], there has not been much elaboration on requirement engineering. As mentioned in the paper, software engineers are generally reluctant to rely on LLMs for higher-level design goals. However, with LLMs achieving remarkable improvements in contextual analysis and reasoning abilities through various methods like prompt engineering and Chain-of-Thought (COT) [16], their applications in requirement engineering are gradually increasing. Table I summarizes and categorizes the tasks in software engineering. Since the collection primarily focuses on the latter half of 2023 and before December 2024, and some papers address multiple tasks, the table reflects the number of tasks after accounting for overlaps.
@@
43
Figure 2: Paper Distribution.
While other works have surveyed LLMs applications in some SE tasks [17, 8, 18], they lack a wider coverage of the general SE area to incorporate recent research developments. More importantly, a focus of LLMs is the main contributions of these works, but there is no distinguish the capabilities between LLMs and LLM-based agents. We summarize the difference between our work and others in Table II, this survey addresses these limitations by distinctly analyzing LLMs and LLM-based agents applications across six SE domains, providing a thorough and up-to-date review. From previous research, it is evident that the performance of LLMs in various applications and tasks heavily depends on the model’s inherent capabilities [10]. More importantly, earlier surveys often present findings from papers spanning in a wide range of publication dates, leading to significant content disparities for LLMs in different SE tasks. For instance, research in requirement engineering was relatively nascent, resulting in sparse content in this area in previous surveys. The recent rise of LLM-based agents, with their enhanced capabilities and autonomy, fills these gaps. By focusing on the latest researches and clearly differentiating between LLMs and LLM-based agents, our survey provides a thorough and in-depth overview of how these technologies are applied and new opportunities they bring to SE.
In summary, we have collected a total of 139 papers directly relevant to this topic, covering the six SE domains mentioned earlier as shown in Figure.1. Our analysis distinguishes between LLM and LLM-based agent contributions, offering a comparative overview and addressing the limitations of previous surveys. Considering the novelty nature of the LLM-based agents field and the lack of standardized benchmarks, this work seeks to offer a detailed review that can guide future research and provide a clearer view of the potential of these technologies in SE.
TABLE II: Comparison between Our Work and the Existing Work for LLM in SE.
@@ II-B Methodology
By following previous research efforts [21, 8, 17], we collected relevant papers from DBLP and arXiv, focusing on publications from the latter half of 2023 to December 2024. DBLP offers comprehensive, curated coverage of computer science venues, particularly in software engineering. We chose arXiv for its open-access nature, which makes the latest research results (including preprints) rapidly available [23, 24]. This ensures we capture the most up-to-date studies, often before they appear in broader databases such as Scopus or Web of Science, thus enabling a timely and inclusive survey of emerging trends in LLM-based research.
II-B1 Database Selection and Rationale
Most other survey papers collect papers from common databases such as Scopus and Web of Science. In contrast, our survey specifically selects DBLP and arXiv due to their distinct advantages:
  • DBLP: offers comprehensive, curated coverage of computer science venues, particularly in software engineering, thus ensuring relevance and high quality of retrieved papers.
  • arXiv: provides open-access preprints, which enables rapid access to the most recent research findings in the field of computer science.
  • IEEE Xplore and ACM: Additional databases such as IEEE Xplore and ACM Digital Library could further complement our search by including more peer-reviewed papers from recognized conferences and journals.
II-B2 Paper Retrieval Process
We applied a systematic filtering process to ensure the relevance and quality of the papers included in this study. Specifically, we conducted a keyword-based filtering process to identify SE-related works. We utilized the keywords listed in Table IV, combining them with Boolean operators to perform precise and comprehensive searches:
  • Within each topic, keywords were joined using the OR operator, e.g., (”Software security” OR ”Vulnerability detection”).
  • To ensure relevance, we combined topic keywords with LLM-related terms using the AND operator, e.g., (”Code generation” AND ”Large Language Models”).
TABLE III: Inclusion and Exclusion Criteria for Paper Selection.
Inclusion criteria
1) The paper explicitly involves the application or study of LLMs.
2) The paper demonstrates relevance to Software Engineering tasks.
3) The paper provides sufficient experimental results and evaluations.
Exclusion criteria
1) Papers with fewer than 7 pages.
2) Papers that focus on general Artificial Intelligence not relate to LLMs.
3) Studies that do not address Software Engineering applications or workflows.
4) Grey literature, such as blog posts, white papers.
5) Duplicate publications.
6) Papers written in languages other than English.
II-B3 Selection Criteria
Following the initial keyword-based retrieval, we applied additional inclusion and exclusion criteria to refine the selected papers further. We manually reviewed titles, abstracts, and, where necessary, the full texts to confirm that each paper specifically addressed the intersection of software engineering and large language models. We clearly established inclusion and exclusion criteria to filter papers, as detailed in Table III. Papers primarily focused on unrelated domains or those mentioning LLMs without substantial relevance to software engineering tasks were excluded. Furthermore, we ensured that selected publications originated from reputable venues, including peer-reviewed journals, conferences, and workshops recognized by the software engineering and artificial intelligence research communities. This rigorous filtering process ensured the final set of papers provided a representative and high-quality overview of recent advancements in applying large language models to software engineering.
Overall, we included papers that explicitly focused on topics involving LLMs, contained keywords from Table IV related to Software Engineering applications, and were at least seven pages long to ensure substantial contributions. We excluded papers unrelated to SE or LLMs, shorter than seven pages, or lacking relevant keywords.
II-B4 Manual Screening and Paper Distribution
After the initial keyword-based retrieval and selection criteria, we also manually reviewed titles, abstracts, and, where necessary, the full texts to confirm that each paper specifically addressed the intersection of software engineering and large language models. Papers primarily focused on unrelated domains or those mentioning LLMs without substantial relevance to software engineering tasks were excluded. Furthermore, we ensured that selected publications originated from reputable venues, including peer-reviewed journals, conferences, and workshops recognized by the software engineering and artificial intelligence research communities. Additionally, we employed a snowballing search technique, systematically reviewing references cited within relevant papers to identify other significant studies that might have been overlooked initially [17]. This filtering process ensured the final set of papers provided a representative and high-quality overview of recent advancements in applying large language models to software engineering.
Overall, we identified 139 relevant papers. Figure 2 presents the distribution of these papers across the six SE domains and the proportion of LLM-based agent studies. However, some papers can be counted as multi-class fields, so the literature review in the figure includes more than 139 total paper instances.
The venue distribution revealed in Figure 3 demonstrates both the interdisciplinary nature and rigorous academic foundation of this research domain. The prominence of arXiv (40.3%) reflects the field’s rapid evolution, where researchers pr ioritize immediate community dissemination of LLM advancements in software engineering. Notably, 59.7% of papers originate from established peer-reviewed venues, with strong representation in software engineering flagships (ICSE: 6.5%, ESEC/FSE: 3.6% combined) and AI/ML top conferences (NeurIPS: 10.1%, ICLR: 2.9%). This dual concentration underscores how LLM-based SE research bridges fundamental AI innovation with practical engineering applications.
The distribution reveals three notable patterns: First, the presence of requirements engineering venues (IEEE RE Conference: 2.9%) alongside code-centric conferences (OOPSLA: 1.4%) illustrates LLMs’ expanding role across the SE lifecycle. Second, NLP venues (ACL/EMNLP: 5.0%) contribute specialized language modeling insights critical for code-text cross-modal tasks. Third, the inclusion of security (ACSAC) [25], HCI (IUI) [26], and software practice (XP) [27] venues demonstrates the field’s maturation into specialized subdomains. This balanced distribution ensures methodological rigor while capturing real-time innovation, validating our survey’s comprehensive view of how LLM advancements redefine software engineering paradigms through contextual adaptation, open-ended reasoning, and human-AI collaboration.
Figure 3: Venue Distribution.
@@ II-C Overall Structure of the Work
The remainder of this paper is organized as follows Section 2 Introduces the architectures and background of LLMs and LLM-based agents, including an overview of RAG, tool utilization, and their implications for SE. Section III-IX is the comparative analysis which Summarizes and compares the datasets, tasks, benchmarks, and metrics used in LLM and LLM-based agents studies across the six SE domains. Section X is the general discussion, Section XI is the final conclusion.
TABLE IV: Keywords for Software Engineering Topics.
@@ III-A Large Language Model
There is an inherent connection between large language models and natural language processing (NLP), with the historical development of natural language technologies tracing back to the 1950s. The earliest attempts to generate language dialogues through machines using specific rules can be traced to the period between 1950 and 1970. The advent of machine learning technologies in the 1980s and the groundbreaking introduction of neural networks in the 1990s indicated a new era for NLP  [28]. These advancements facilitated significant progress in the NLP field, especially in the development of technologies for text translation and generation. The development of Long Short-Term Memory (LSTM) and Recurrent Neural Networks (RNN) during this period enabled more effective handling of the sequential nature of language data [29, 30]. These models addressed challenges associated with the lack of dependency in context, thereby enhancing the application of NLP in various domains.
@@ III-C Large Language Model Based Agent
Figure 4: Illustration of Common Data Augmentation Methods.
Nowadays, various agent systems are emerging and they rely on large language models to make judgments, combined with techniques such as few-shot learning and multi-turn dialogue for model fine-tuning. However, due to the lack of datasets many researchers employ different methods for data augmentation. Common approaches include synonym replacement, where words in the text are replaced with synonyms from the same domain to increase textual diversity; back-translation, where the text is translated into another language and then back to the original language to generate new texts with slightly different grammatical structures and word choices; Paraphrasing refers to the new dialogue which similar in context but slightly different in expression created through manual or automated means; Synthetic Data Generation refers to use pretrained model to make a synthetic data generation as shown in Figure.4. In 2023, Chenxi Whitehouse explored using LLMs for data augmentation to enhance the performance of multilingual commonsense reasoning datasets, especially under conditions of extremely limited training data [44]. The study employed various LLMs (such as Dolly-v2, StableVicuna, ChatGPT, and GPT-4) to generate new data. These models were prompted to create new examples similar to the original data, thereby increasing the diversity and quantity of training data. Prompt engineering was mentioned as an essential skill for effectively interacting with LLMs. By applying these prompt patterns, users can efficiently customize their dialogues with LLMs ensuring the generation of high-quality outputs and achieving complex automated tasks. In 2023, Jules White introduced a set of methods and patterns to enhance prompt engineering, optimizing interactions with LLMs such as ChatGPT [45]. This study categorizes prompt engineering into five main areas: Input Semantics, Output Customization, Error Identification, Prompt Improvement, and Interaction, to address a wide range of problems and adapt to different fields.
@@ III-D Single and Multi-Agents
Multi-agent systems involve the collaboration of multiple LLMs or agents to tackle complex tasks effectively. These systems fully utilize the advantages of multiple models, with each model specializing in specific aspects of the task to reduce overhead caused by multi-processes in single agents, the collaboration among agents allows for more sophisticated and robust problem-solving capabilities. Due to their exceptional capabilities, more researchers are beginning to explore the field of Multi-LLM based agents and start applying into software engineering domains. In 2024, a lot of researchers adopt the multi-agent system into the practical experiments [48, 49].
@@ III-E LLM in Software Engineering
Currently, there is no comprehensive and widely accepted definition specifying the exact capabilities an LLM must exhibit to be considered an LLM-based agent, particularly within the context of software engineering. Given the extensive application of LLMs in software engineering and the varying levels of autonomy and intelligence exhibited by current frameworks, it is critical to establish a clear distinction between general LLM frameworks and those that can be defined as agents. This study synthesizes criteria from mainstream definitions and literature from the first half of 2024, reflecting the consensus found across multiple prominent research contributions.
Specifically, our criteria are derived from recent works emphasizing autonomous reasoning, decision-making, and adaptive planning capabilities in LLM-based systems, such as those outlined in frameworks including SWE-agent [53], CODEAGENT [54], AutoCodeRover [55], and GOEX [56]. Further foundational concepts from broader agent literature, such as MetaGPT [57], AgentVerse [58], Reflexion [59], and ExpeL [40], support these definitions by consistently demonstrating advanced agent capabilities.
Based on this extensive review, we define an LLM architecture as an agent if it meets most of the criteria listed in Table V. Criteria 1 to 4 (LLM as central reasoning core, decision-making and planning, autonomous tool usage, and evaluation capabilities) are considered fundamental for a basic level of agency. Criteria 5 and 6 (contextual multi-turn interaction and autonomous learning capability) indicate more advanced agentic behaviors, beneficial but not strictly necessary for an LLM to be initially classified as an agent. This distinction provides flexibility while preserving rigor, allowing diverse frameworks to be effectively and consistently evaluated.
TABLE V: Criteria for LLM-based agent
@@ IV-A LLMs Tasks
Requirement Classification and Extraction. In the field of requirement engineering, LLMs have demonstrated significant potential in automating and enhancing tasks such as requirement elicitation, classification, generation, specification generation, and quality assessment. Requirement classification and extraction is a crucial task in requirement engineering during the development process. It is common to encounter situations where clients present multiple requirements at once, necessitating manual classification by developers. By categorizing requirements into functional and non-functional requirements, developer can better understand and manage them, thanks to the strong performance of LLMs in classification tasks, many relevant frameworks have been developed. The PRCBERT framework, utilizing the BERT pre-trained language model, transforms classification problems into a series of binary classification tasks through flexible prompt templates, significantly improving classification performance [60]. Studies have shown that the PRCBERT achieved an F1 score of 96.13% on the PROMISE dataset which outperform the previous state-of-arts NoRBERT [61] and BERT-MLM models [36]. Additionally, the application of ChatGPT in requirement information retrieval has shown promising results, by classifying and extracting information from requirement documents, ChatGPT achieved comparable or even better $F\beta$ scores under zero-shot settings, particularly in feature extraction tasks, where its performance surpassed baseline models [62]. A comparative study further evaluated ChatGPT (GPT-3.5 and GPT-4) against traditional models such as SVM and LSTM across five datasets—PROMISE, Dronology, ReqView, Leeds Library, and WASP. Using a class-specific $F\beta$ scoring method to emphasize recall, the results demonstrated that GPT-3.5 Few-Shot prompting achieved the best overall classification performance, while GPT-4 Zero-Shot showed notable strength in specialized categories like OnlyFunctional requirements [63]. As seen in Table I, there is also substantial literature and research on using LLMs to automatically generate requirements and descriptions in requirement engineering.
Requirement Elicitation and Specification Generation. By automating the generation and description of requirements, the efficiency and accuracy of requirement elicitation can be improved. Research indicates that LLMs hold significant potential in requirements generation task. For example, using ChatGPT to generate and gather user requirements, studies found that participants with professional knowledge could use ChatGPT more effectively, indicating the influence of domain expertise on the effectiveness of LLM-assisted requirement elicitation [64]. The study employed qualitative assessments of the LLMs’ output against predefined criteria for requirements matches, including full matches, partial matches, and the relevancy of the elicited requirements, although their success varied depending on the complexity of the task and the experience of the users, the result showing that LLMs could effectively assist in eliciting requirements, and its particularly useful in identifying, and suggesting requirements based on the large corpus of training data they provided. The GeneUS pipeline [65] pushes this boundary by unifying user story and test case generation through a three-stage prompting process. Evaluated using the RUST framework (readability, understandability, specifiability, technical aspects), GPT-4-generated artifacts received high ratings, particularly in understandability and specifiability, showing LLMs’ potential in supporting agile requirement processes.
@@
The SRS (Software Requirement Specification) generation is an important task which the developer normally spent a lot of time to refine and verified. In [66], researchers use both iterative prompting and a single comprehensive prompt to assess the performance of LLMs to generate SRS. The experiment conducted on GPT-4 and CodeLlama-34b one close-source LLM and one open-source LLM for comprehensive evaluation, the generated SRS will compare with human-crafted SRS and finally scored by the likert scale. The result indicate that, the human-generated SRS was overall superior, but CodeLlama often came close, sometimes outperforming in specific categories. The CodeLlama scored higher in completeness and internal consistency than GPT-4 but less concise, so this stuy demonstrated the potential of using fine-tuned LLMs to generate SRS and increase the overall project productivity. Another paper also explores using LLMs for generating specifications. In [67], the authors introduce a framework called SpecGen for generating program specifications. The framework primarily uses GPT-3.5-turbo as the base model and employs prompt engineering combined with multi-turn dialogues to generate the specifications. SpecGen applies four mutation operators to modify these specifications and finally uses a heuristic selection strategy to choose the optimal variant. The results show that SpecGen can generate 70% of the program specifications, outperforming traditional tools like Houdini [68] and Daikon [^1] https://github.com/codespecs/daikon. A significant advancement in formal modeling is shown in PathOCL [69], which enhances OCL constraint generation from UML diagrams by injecting contextual navigation paths into GPT-4 prompts. Evaluated on 15 UML models and 168 English specifications, PathOCL significantly outperformed UML-Augmented prompting on Validity@K and Correctness@K metrics, while reducing hallucinations and improving alignment with formal intent.
@@
Prompt Engineering and Requirement Completeness. Furthermore, designing prompt patterns can significantly enhance LLMs’ capabilities in tasks such as requirement elicitation and system design. The paper provides a catalog of 13 prompt patterns, each aimed at addressing specific challenges in software development [70]. The experiments test the efficacy of these patterns in real world scenarios to validate their usefulness. By applying different prompt patterns, the study found that these patterns could help generate more structured and modular results and reduce common errors. Automated requirement completeness enhancement is another important benefit brought by the LLMs in requirement generation. The study [71] use BERT’s Masked Language Model (MLM) can detect and fill in missing parts in natural language requirements, significantly improving the completeness of requirements. BERT’s MLM achieved a precision of 82%, indicating that 82% of the predicted missing terms were correct.